OpenStudy (anonymous):

valuate the integral (4ln5x-(2/2^(2)-9) my attempt landed me with x(12ln(x)-X^(2)+3(4ln(5)+5))/3

OpenStudy (anonymous):

\[\int\limits \left(\frac{17}{2}+4 \text{ Log}[ 5 x ]\right) \, dx=\frac{9 x}{2}+4 x \log (5 x)+c \]